The Allied Corporation analyzes a project that requires an immediate investment of $445. Allied estimates that at the end of the first year the project will generate a cash flow of $640, but that at the end of the second year, when the project ends, it will generate a negative cash flow of $80. The project's required rate of return is estimated to be 8.75%. Calculate the NPV of Allied's project. $___?___
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=640/1.0875

=$588.51

Present value of outflows=445+(80/1.0875^2)

=445+(80*0.845554234)

=$512.64

NPV=Present value of inflows-Present value of outflows

=$588.51-$512.64

which is equal to

=$75.87(Approx).